OpenStudy (anonymous):

how to solve 27=1/2x-4x-8

OpenStudy (gw2011):

(1/2)(x)-4x-8=27 Multiply each term by 2 to eliminate the denominator and you get: x-8x-16=54 -7x=70 x=-10